# TripNet: Learning Large-scale High-fidelity 3D Car Aerodynamics with Triplane Networks

### v1

@@ C.1 Wall Shear Stress Field
DrivAerNet++ RegDGCNN (Elrefaie et al., 2024a) 13.82 3.64 11.54 36.42 36.57
FigConvNet (Choy et al., 2025) 9.86 2.22 7.84 22.32 17.65
TripNet (Ours) 9.52 2.15 8.14 22.07 17.18
@@ C.2 Pressure Field
Figures A3 and A4 provide a detailed comparison of pressure predictions and errors across multiple deep learning models for three car designs selected from the unseen test set. These figures highlight differences between model predictions and the ground truth CFD results, visualizing the absolute error distribution for each method. We consider various car categories and configurations, including different underbody and wheel setups, to assess the generalization capabilities of deep learning-based aerodynamic prediction methods across diverse design variations.
(a) Car design F_D_WM_WW_1769 (fastback with detailed underbody, with wheels, and with mirrors).
@@
(b) Car design F_S_WWC_WM_504 (fastback with smooth underbody, wheels closed, and with mirrors).
@@
(c) Car design E_S_WWC_WM_094 (estateback with smooth underbody, wheels closed, and with mirrors).
@@
Figure A3: Comparison of pressure predictions and errors for three car designs from the unseen test set. For each car, the first row shows the input mesh followed by predictions from RegDGCNN, FigConvNet, TripNet (ours), and the ground-truth CFD. The second row highlights the absolute difference between the predictions and the ground-truth CFD, visualizing the absolute error distribution for each model.
(a) Car design F_D_WM_WW_1769 (fastback with detailed underbody, with wheels, and with mirrors).
@@
(b) Car design F_S_WWC_WM_504 (fastback with smooth underbody, wheels closed, and with mirrors).
@@
(c) Car design E_S_WWC_WM_094 (estateback with smooth underbody, wheels closed, and with mirrors).
@@
Figure A4: Comparison of pressure predictions and errors for three car designs from the unseen test set. For each car, the first row shows the input mesh followed by predictions from RegDGCNN, FigConvNet, TripNet (ours), and the ground-truth CFD. The second row highlights the absolute difference between the predictions and the ground-truth CFD, visualizing the absolute error distribution for each model.
(a) Car design F_D_WM_WW_1769 (fastback with detailed underbody, with wheels, and with mirrors).
@@
(b) Car design F_S_WWC_WM_504 (fastback with smooth underbody, wheels closed, and with mirrors).
@@
(c) Car design E_S_WWC_WM_094 (estateback with smooth underbody, wheels closed, and with mirrors).
@@
Figure A5: Comparison of wall shear stress predictions and errors for three car designs from the unseen test set. For each car, the first row shows the input mesh followed by predictions from RegDGCNN, FigConvNet, TripNet (ours), and the ground-truth CFD. The second row highlights the absolute difference between the predictions and the ground-truth CFD, visualizing the absolute error distribution for each model.
(a) Car design F_D_WM_WW_1769 (fastback with detailed underbody, with wheels, and with mirrors).
@@
(b) Car design F_S_WWC_WM_504 (fastback with smooth underbody, wheels closed, and with mirrors).
@@
(c) Car design E_S_WWC_WM_094 (estateback with smooth underbody, wheels closed, and with mirrors).
@@
Figure A6: Comparison of wall shear stress predictions and errors for three car designs from the unseen test set. For each car, the first row shows the input mesh followed by predictions from RegDGCNN, FigConvNet, TripNet (ours), and the ground-truth CFD. The second row highlights the absolute difference between the predictions and the ground-truth CFD, visualizing the absolute error distribution for each model.
@@ Appendix D Turbulence Modeling
The derivations and turbulence model discussed in this section are summarized from (Wilcox et al., 1998; Menter et al., 2003, 2021; Menter, 1993; Spalding, 1974), providing a foundation for advanced turbulence modeling approaches capable of addressing the challenges of accurately simulating real-world turbulent flows.
@@
Reynolds averaging is introduced as a key approach to address the complexity of turbulent flows by decomposing the instantaneous flow variables into their mean and fluctuating components. However, this approach leads to the well-known ”closure problem” due to the additional unknown terms introduced in the Reynolds-averaged equations. To resolve this, turbulence models are required to provide approximations for the unknown terms. These models, employed within the Reynolds-Averaged Navier-Stokes (RANS) framework, enable practical solutions for complex turbulent flows.
@@
The $k-\omega$ Shear Stress Transport (SST) model is a widely used two-equation turbulence model that solves for turbulence kinetic energy ($k$) and the specific dissipation rate ($\omega$). The SST model is particularly effective in capturing flow separation, which is critical for accurately predicting aerodynamic and hydrodynamic performance. Both the DrivAerNet (Elrefaie et al., 2024a) and DrivAerNet++ (Elrefaie et al., 2024d) datasets are generated using RANS simulations with the $k-\omega$ SST turbulence model, ensuring high-fidelity aerodynamic predictions.
@@ Appendix E Description of the DrivAerNet++ Dataset
In this study, we utilize the DrivAerNet++ dataset, introduced by Elrefaie et. al (Elrefaie et al., 2024a, d), a comprehensive multimodal dataset specifically developed for aerodynamic car design. It includes over 8,000 unique industry-standard car models, each meticulously modeled using high-fidelity CFD simulations. The dataset features a variety of car body styles, such as fastback, notchback, and estateback, and encompasses various underbody and wheel configurations applicable to both internal combustion and electric cars.
Table A4: Comparison of drag coefficient ($C_{d}$) ranges and mean values across different car configurations in the DrivAerNet++ dataset. Each row represents a specific configuration with the corresponding number of samples, $C_{d}$ range, and mean $C_{d}$. Each car is normalized by its corresponding projected frontal area. The images illustrate the respective car designs used for the aerodynamic simulations. Two underbody configurations are considered: smooth and detailed, and three wheels/tires configurations: WW (with wheels), WWS (smooth), and WWC (with wheels closed).
@@
Accompanying each model are detailed 3D meshes, parametric models, aerodynamic coefficients, as well as extensive flow and surface field data. DrivAerNet++ also provides segmented parts for car classification and point cloud data, enabling a broad spectrum of machine learning applications including design optimization, generative modeling, surrogate model training, CFD simulation acceleration, and geometric classification. This dataset represents a significant advancement over existing datasets due to its size, realistic industry-standard car shapes, shape diversity, and higher fidelity of its CFD simulations, making it an invaluable resource for training deep learning models. Table A4 provides a detailed comparison of drag coefficient ($C_{d}$) ranges, mean values, and sample sizes across various DrivAerNet++ car configurations.

### v2

@@ B.2 Model Performance
Transolver 1024 7.72 1.56 3.00 28.35 26.01
4096 5.60 1.35 3.44 23.82 22.54
16384 5.58 1.29 5.03 22.72 21.49
32768 5.37 1.28 5.89 22.52 21.31
FigConvNet 1024 11.84 2.00 3.47 35.06 33.36
4096 5.66 1.29 3.33 22.59 21.37
16384 4.77 1.14 4.88 20.50 18.89
32768 4.38 1.13 5.73 20.98 18.59
TripNet 1024 4.11 1.15 3.23 20.87 18.67
4096 4.18 1.17 3.41 20.77 18.93
16384 4.08 1.13 5.45 20.22 18.43
32768 4.23 1.11 5.52 20.35 18.52
Appendix C Additional Results
C.1 Drag coefficient
Figure 3 illustrates the architecture of a convolutional neural network designed to predict the drag coefficient ($C_{d}$) from stacked triplanes of 3D geometry. This surrogate model follows an encoder-style structure, progressively reducing spatial dimensions while increasing feature depth, followed by fully connected layers for regression. The model begins with an input tensor of shape $9\times 128^{2}$, representing 9-channel triplanes $\mathbf{\tilde{P}}_{xy}$, $\mathbf{\tilde{P}}_{yz}$ and $\mathbf{\tilde{P}}_{xz}$. The architecture consists of convolution encoder and flattening dense layers. The output layer predicts a scalar drag coefficient $C_{d}$. Table  5 summarizes the performance of different models on drag coefficient prediction on the DrivAerNet++ dataset. TripNet achieves the state-of-the-art performance with $R^{2}$ score of 0.957.
Figure 3: Lightweight CNN model for drag coefficient prediction
Table 5: Performance comparison on drag coefficient prediction DrivAerNet++ test set.
(×$(\times 10^{-5})$ (×$(\times 10^{-3})$ (×$(\times 10^{-2})$
DrivAerNet++ GCNN Kipf and Welling [2016] * 17.1 10.43 15.03 0.596
RegDGCNN Elrefaie et al. [2024b]* 14.2 9.31 12.79 0.641
PointNet Qi et al. [2017a]* 14.9 9.60 12.45 0.643
TripNet (Ours) 9.1 7.17 7.70 0.957
@@ C.2 Pressure Field
Figures 4 and 5 provide a detailed comparison of pressure predictions and errors across multiple deep learning models for three car designs selected from the unseen test set. These figures highlight differences between model predictions and the ground truth CFD results, visualizing the absolute error distribution for each method. We consider various car categories and configurations, including different underbody and wheel setups, to assess the generalization capabilities of deep learning-based aerodynamic prediction methods across diverse design variations.
(a) Design F_D_WM_WW_1769 (fastback, detailed underbody, with wheels, and with mirrors).
@@
(b) Design F_S_WWC_WM_504 (fastback, smooth underbody, wheels closed, and with mirrors).
@@
(c) Design E_S_WWC_WM_094 (estateback, smooth underbody, wheels closed, and with mirrors).
(d) Design N_S_WW_WM_025 (notchback, smooth underbody, with wheels, and with mirrors).
@@
Figure 4: Comparison of pressure predictions and errors for three car designs from the unseen test set. For each car, the first row shows the input mesh followed by predictions from RegDGCNN, FigConvNet, Transolver, TripNet (ours), and the ground-truth CFD. The second row highlights the absolute difference between the predictions and the ground-truth CFD, visualizing the absolute error distribution for each model.
(a) Design F_D_WM_WW_1769 (fastback, detailed underbody, with wheels, and with mirrors).
@@
(b) Design F_S_WWC_WM_504 (fastback, smooth underbody, wheels closed, and with mirrors).
@@
(c) Design E_S_WWC_WM_094 (estateback, smooth underbody, wheels closed, and with mirrors).
(d) Design N_S_WW_WM_025 (notchback, smooth underbody, with wheels, and with mirrors).
@@
Figure 5: Comparison of pressure predictions and errors for three car designs from the unseen test set. For each car, the first row shows the input mesh followed by predictions from RegDGCNN, FigConvNet, TripNet (ours), and the ground-truth CFD. The second row highlights the absolute difference between the predictions and the ground-truth CFD, visualizing the absolute error distribution for each model.
C.3 Wall Shear Stress Field
Figures 6 and 7 provide a detailed comparison of wall shear stress predictions and errors across multiple deep learning models for three car designs selected from the unseen test set. These figures highlight differences between model predictions and the ground truth CFD results, visualizing the absolute error distribution for each method. We consider various car categories and configurations, including different underbody and wheel setups, to assess the generalization capabilities of deep learning-based aerodynamic prediction methods across diverse design variations.
(a) Design F_D_WM_WW_1769 (fastback, detailed underbody, with wheels, and with mirrors).
@@ C.3 Wall Shear Stress Field
(b) Design F_S_WWC_WM_504 (fastback, smooth underbody, wheels closed, and with mirrors).
@@
(c) Design E_S_WWC_WM_094 (estateback, smooth underbody, wheels closed, and with mirrors).
(d) Design N_S_WW_WM_025 (notchback, smooth underbody, with wheels, and with mirrors).
@@
Figure 6: Comparison of wall shear stress predictions and errors for three car designs from the unseen test set. For each car, the first row shows the input mesh followed by predictions from RegDGCNN, FigConvNet, Transolver, TripNet (ours), and the ground-truth CFD. The second row highlights the absolute difference between the predictions and the ground-truth CFD, visualizing the absolute error distribution for each model.
(a) Design F_D_WM_WW_1769 (fastback, detailed underbody, with wheels, and with mirrors).
@@
(b) Design F_S_WWC_WM_504 (fastback, smooth underbody, wheels closed, and with mirrors).
@@
(c) Design E_S_WWC_WM_094 (estateback, smooth underbody, wheels closed, and with mirrors).
(d) Design N_S_WW_WM_025 (notchback, smooth underbody, with wheels, and with mirrors).
@@
Figure 7: Comparison of wall shear stress predictions and errors for three car designs from the unseen test set. For each car, the first row shows the input mesh followed by predictions from RegDGCNN, FigConvNet, Transolver, TripNet (ours), and the ground-truth CFD. The second row highlights the absolute difference between the predictions and the ground-truth CFD, visualizing the absolute error distribution for each model.
C.4 Volumetric Flow Field
As shown in Figure  8, TripNet predictions closely match the CFD ground truth, capturing key flow features near the vehicle surface and in the wake region. The high level of agreement across upstream and downstream sections demonstrates TripNet’s spatial accuracy and generalization capabilities.
Figure 8: Cross-sectional comparison of volumetric field predictions between CFD (left) and TripNet (right) at multiple locations along the car body
@@ Appendix D Turbulence Modeling
The derivations and turbulence model discussed in this section are summarized from Wilcox et al. [1998], Menter et al. [2003, 2021], Menter [1993], Spalding [1974], providing a foundation for advanced turbulence modeling approaches capable of addressing the challenges of accurately simulating real-world turbulent flows.
@@
Reynolds averaging is introduced as a key approach to address the complexity of turbulent flows by decomposing the instantaneous flow variables into their mean and fluctuating components. However, this approach leads to the well-known "closure problem" due to the additional unknown terms introduced in the Reynolds-averaged equations. To resolve this, turbulence models are required to provide approximations for the unknown terms. These models, employed within the Reynolds-Averaged Navier-Stokes (RANS) framework, enable practical solutions for complex turbulent flows.
@@
The $k-\omega$ Shear Stress Transport (SST) model is a widely used two-equation turbulence model that solves for turbulence kinetic energy ($k$) and the specific dissipation rate ($\omega$). The SST model is particularly effective in capturing flow separation, which is critical for accurately predicting aerodynamic and hydrodynamic performance. Both the DrivAerNet Elrefaie et al. [2024b] and DrivAerNet++ Elrefaie et al. [2024d] datasets are generated using RANS simulations with the $k-\omega$ SST turbulence model, ensuring high-fidelity aerodynamic predictions.
@@ Appendix E Description of the DrivAerNet++ Dataset
In this study, we utilize the DrivAerNet++ dataset, introduced by Elrefaie et al. [2024b, d], a comprehensive multimodal dataset specifically developed for aerodynamic car design. It includes over 8,000 unique industry-standard car models, each meticulously modeled using high-fidelity CFD simulations. The dataset features a variety of car body styles, such as fastback, notchback, and estateback, and encompasses various underbody and wheel configurations applicable to both internal combustion and electric cars.
Table 6: Comparison of drag coefficient ($C_{d}$) ranges and mean values across different car configurations in the DrivAerNet++ dataset. Each row represents a specific configuration with the corresponding number of samples, $C_{d}$ range, and mean $C_{d}$. Each car is normalized by its corresponding projected frontal area. The images illustrate the respective car designs used for the aerodynamic simulations. Two underbody configurations are considered: smooth and detailed, and three wheels/tires configurations: WW (with wheels), WWS (smooth), and WWC (with wheels closed).
@@
Accompanying each model are detailed 3D meshes, parametric models, aerodynamic coefficients, as well as extensive flow and surface field data. DrivAerNet++ also provides segmented parts for car classification and point cloud data, enabling a broad spectrum of machine learning applications including design optimization, generative modeling, surrogate model training, CFD simulation acceleration, and geometric classification. This dataset represents a significant advancement over existing datasets due to its size, realistic industry-standard car shapes, shape diversity, and higher fidelity of its CFD simulations, making it an invaluable resource for training deep learning models. Table 6 provides a detailed comparison of drag coefficient ($C_{d}$) ranges, mean values, and sample sizes across various DrivAerNet++ car configurations.